Please me with this equation Simp[lify sqrt(75a^2) My answer is 0 is this corrrect. thanks for all that you do.
Answer by jai_kos(139)

√ (75a^2)
Solution:
√75a^2 = √ 3* 25 * a^2
√75a^2 = √ 3* 5^2 * a^2
√75a^2 = 5 * a √ 3
√75a^2 = 5a√ 3